pliopikk:

Nikhil gets paid a 5 percent commission on every pair of shoes that he sells. He earned $1.00 on the last pair of shoes that he sold. The expression that can be used to represent x, the price of the shoes, is 0.05x=1 What was the price of the shoes?

dude:

I'm assuming they gave you the equation 0.05x=1 You want to solve for x, so divide by 0.5 on both sides \(x=\large\frac{1}{0.05}\)

pliopikk:

what the answer fool

dude:

I am not allowed to give you a direct answer but I just showed you how to get to it. \(\bf \color{red}{x=\large\frac{1}{0.05}}\)

pliopikk:

what is x

dude:

x is the cost per shoe (They gave you this in the equation)

dude:

Question*

pliopikk:

why not just a answer u going to get in trouble are somthing

dude:

Rules: https://questioncove.com/rules

Afterdeath:

@pliopikk, we're not allowed to give u the answer so x=1/0.05, try that and then maybe, just maybe you'll hv ur answer

pliopikk:

wth is that

pliopikk:

5.00 is the answer

dude:

No Divide 1 by 0.05 That is what \(\large \frac{1}{0.05}\) means

pliopikk:

it 20.00

dude:

Good job

pliopikk:

thank u